# SoK: Liquid Staking Tokens (LSTs) and Emerging Trends in Restaking

### v2

@@ III-E Liquid Staking Tokens in DeFi
LSTs can be freely transferred and used in various DeFi applications. Reward-based LSTs are more popular in DeFi because rebase tokens are not widely compatible with most protocols. This section explores the main uses of reward-based LSTs in DeFi and explains why rebase tokens are less suited for these use-cases.
@@
LSTs can be traded on DEXs and CEXs at market prices tah may differ from the protocol prices. Most AMM-based DEX platforms, such as Uniswap, do not support trading of rebase LSTs [33]. However, rebase LSTs can be traded on Curve v2 [33].
@@
LSTs can be used as collateral in DeFi lending protocols to borrow other tokens. Reward-based LSTs are particularly favored as their value appreciates over time due to staking rewards, making them a dominant collateral asset in DeFi, surpassing ETH [8]. Rebase-LSTs, however, are not used as collateral because staking rewards would need to be divided between the borrower and the lender. In contrast, for reward-LSTs, all staking rewards accrue to the collateral provider.
@@
LSTs are also used as collateral for minting decentralized stablecoins. The appreciating value of LSTs, driven by staking rewards, reduces the liquidation risk of the collateral over time, enhancing stability and efficiency in stablecoin systems.
Leveraging
LSTs can be leveraged in lending protocols through a process known as looping. This involves using LSTs as collateral to borrow ETH, which is then used to purchase additional LSTs. Another leveraging method involves flash loans, where a flash loan is used to purchase LSTs, lock them as collateral in a lending protocol, borrow ETH, and repay the loan.
Providing Liquidity to AMM-DEX
LSTs can be utilized in liquidity provision to AMM-based DEXs. In LST-ETH liquidity pools, part of the LSTs is converted to ETH, meaning only a fraction of the liquidity pool earns staking rewards. High trading volumes and corresponding swap fees are necessary to compensate for missed staking opportunities. Since future LST prices can be estimated based on the current staking rate, LSTs are frequently traded in AMMs with concentrated liquidity models, such as Uniswap v3.
@@
Operating an Ethereum validator typically requires a deposit of 32 ETH. However, with LSPs, it is possible to operate a validator with only 4 ETH or 8 ETH, with the remaining ETH provided by the staking pool of the LSP.
The risks inherent in LST-based DeFi, particularly in scenarios where LST de-pegs result in liquidations, are becoming increasingly prominent in the literature and are discussed in Section VII.
@@ IV Comparison of Liquid Staking Protocols
Liquid staking is the largest category among DeFi protocols by TVL and the preferred method of staking for users. Currently, 37% of all staked ETH is held via Liquid Staking Tokens (LSTs). Table III lists the major Liquid Staking Protocols (LSPs), their total value locked (TVL), market share within the broader ETH staking market, and other relevant metrics.
Table III: Overview of major Liquid Staking Protocols (LSPs), highlighting total value locked (TVL), market share of total staked ETH, and associated metrics.
Protocol Reference TVL (\$B) Market (%) LSTs Governance Token Nodes Validators Fees (%)
@@ IV-A Major Liquid Staking Protocols
This section discusses the architecture of major Liquid Staking Protocols (LSPs). The selected protocols were chosen based on the value of staked ETH [10] and represent all previously described mechanisms. A summary of the design decisions for these LSPs is presented in Table IV.
Table IV: Taxonomy of Liquid Staking Protocols (LSPs) comparing permissionless access and collateral requirements to operate validator and visibility of validator sets.
@@ IV-B Empirical Analysis
Table V presents the taxonomy of the major LSTs that were empirically analyzed in this study, with most tokens operating in the reward-based model and naively launched on Ethereum. The differentiations between rebase and reward models for staking reward distribution is crucial for correct empirical analysis of the tokens.
Table V: Token taxonomy for Liquid Staking Tokens (LSTs) with staking reward distribution models supported blockchains. The analyzed tokens are naively minted only on Ethereum and bridged to other blockchain.
@@
StakeWise sETH2, rETH2 Dual Ethereum
@@
The dispersion analysis of ETH LSTs is presented in Figure 7 and supplementary Figure 9 in Appendix. This analysis highlights: (i) The distribution of daily return differences between reward-LSTs and daily staking rewards. ii) The deviation from the target value of 1 ETH for rebase-LSTs.
@@ V-E Comparison of Restaking Protocols
Restaking and liquid restaking represent emerging advancements in the DeFi space, garnering significant user attention and capital. Collectively, restaking protocols boast a TVL of over \$25.2 billion, with the majority concentrated in three major protocols: EigenLayer (\$15.07 billion), Babylon Protocol (\$5.5 billion), and Symbiotic (\$2.2 billion)[42]. Notably, EigenLayer is the third-largest DeFi protocol by TVL. A brief overview of these protocols is given in Table VI.
Table VI: Overview of major restaking protocols, highlighting total value locked (TVL), underlying assets, Actively Validated Services (AVS), and associated metrics.
Protocol Reference TVL (\$B) Underlying Assets Restaking Rewards Governance Token AVS

### v3

@@ III-E Liquid Staking Tokens in DeFi
Decentralized Finance (DeFi), blockchain-based financial services [35, 7, 36], offers an extensive array of services, including trading, lending, and digital asset management [37]. The advantage of LSTs over native staking is the fact that they can be freely transferred and used in various DeFi protocols. Reward-based LSTs are more popular in DeFi because rebase tokens are not widely compatible with most protocols, as explained below. This section further presents the main uses of reward-based LSTs in DeFi.
@@
Automated Market Markers (AMMs) that are the backbone of DEX allow and trade, list or provide liquidity to any token pairs [38]. LSTs can be traded on DEXs and CEXs at market prices at any time without exit queues that are an inherent component of native staking. However, the market price of LST on the exchange may differ from the protocol prices. This price dispcrepancies might lead to the arbitrage opportunities, presented earlier in this chapter. Most AMM-based DEXs, including the largest DEX - Uniswap, do not support trading of rebase LSTs [33]. These tokens can be traded on Curve DEX [33]. Table III depicts how LSTs should be listed at AMM-DEX in order to minimize trading fees and optimize capital efficiency [39].
TABLE III: Optimal AMMs for Trading Pairs with Various LST Types Based on Capital Efficiency. Concentrated Liquidity Market Maker (CLMM) [40] Requires Periodic Rebalancing [39].
Token 1 Token 2 AMMs
rebase-LST ETH Stableswap [41]
rebase-LST rebase-LST Stableswap
reward-LST ETH Cryptoswap [42], CLMM(*)
reward-LST reward-LST Cryptoswap, CLMM(*)
@@
LSTs can be used as collateral in DeFi lending protocols to borrow other tokens [43]. Reward-based LSTs are particularly favored as their value appreciates over time due to staking rewards, making them a dominant collateral asset in DeFi, surpassing ETH [8]. Rebase-LSTs, however, are not used as collateral because staking rewards would need to be divided between the borrower and the lender. In contrast, for reward-LSTs, all staking rewards accrue to the collateral provider.
Leveraging
Algorithm 5 Leveraging Strategy for LSTs
1:  Deposit LSTs as collateral in a lending protocol.
2:  Borrow ETH against the collateralized LSTs.
3:  Buy additional LSTs using the borrowed ETH on a DEX.
4:  Repeat steps 1-3 to increase leverage, ensuring the collateralization ratio remains above the protocol’s liquidation threshold.
LSTs can be leveraged in lending protocols [44, 45, 46] through a process known as looping [47]. This strategy, depicted in Algorithm 5, involves using LSTs as collateral to borrow ETH, which is then used to purchase additional LSTs.
Algorithm 6 Leveraging LSTs Strategy with Flash Loan
1:  Borrow a flash loan in ETH from a lending protocol.
2:  Buy LSTs with the borrowed ETH on a DEX.
3:  Deposit the purchased LSTs as collateral in a lending protocol.
4:  Borrow ETH against the collateralized LSTs.
5:  Repay the flash loan using the borrowed ETH.
The more capital-efficient leveraging method applies flash loans. In this strategy, presented in Algorithm 6 a flash loan is used to purchase LSTs, lock them as collateral in a lending protocol, borrow ETH, and repay the loan. According to emirical research, most LST leveraging strategies gained higher returns compared to native staking [48]. However, the DeFi leverage is exposed to liquidation risks [44, 45, 46].
Providing Liquidity to AMM-DEX
LSTs can be utilized in liquidity provision to AMM-based DEXs [49, 50]. In LST-ETH liquidity pools, part of the LSTs is converted to ETH, meaning only a fraction of the liquidity pool earns staking rewards. High trading volumes and corresponding swap fees are necessary to compensate for missed staking opportunities. Since future LST prices can be estimated based on the current staking rate, LSTs are frequently traded in AMMs with concentrated liquidity models, such as Uniswap v3. However, provided liquidity to AMM pool with LSTs does not always provide higher rewards than native staking due to impermanent loss [49, 50]. The topic of profitability of various DeFi strategies with LSTs is further discussed in Section VII, as empirical research showed that most of LP-strategies involving LSTs leads to the lower returns than native staking [39, 48].
@@
Decentralized stablecoins are DeFi tokens that maintain their value pegged to a fiat currency or commodity thanks to the on-chain assets that backed them [51, 52]. LSTs, as well as LP-positions in AMMs including LSTs, can be provided as collateral to mint decentralized stablecoins[53, 54]. The appreciating value of LSTs, driven by staking rewards, reduces the liquidation risk of the collateral over time, enhancing stability and efficiency in stablecoin systems.
@@
Operating an Ethereum validator typically requires a deposit of 32 ETH. However, with LSPs, it is possible to operate a validator with only 4 ETH [55] or 8 ETH [19], with the remaining ETH provided by the staking pool of the LSP.
The risks inherent in LST-based DeFi, particularly in scenarios where LST de-pegs result in liquidations [56], are becoming increasingly prominent in the literature and are discussed in Section VII.
@@ IV Comparison of Liquid Staking Protocols
Liquid staking is the largest category among DeFi protocols by TVL and the preferred method of staking for users. Currently, 37% of all staked ETH is held via Liquid Staking Tokens (LSTs). Table IV lists the major Liquid Staking Protocols (LSPs), their total value locked (TVL), market share within the broader ETH staking market, and other relevant metrics.
TABLE IV: Overview of major Liquid Staking Protocols (LSPs), highlighting total value locked (TVL), market share of total staked ETH, governance (Gov.) token and associated metrics.
Protocol TVL (\$B) Market (%) LSTs Gov. Token Nodes Validators Fees (%)
@@ IV-A Major Liquid Staking Protocols
This section discusses the architecture of major Liquid Staking Protocols (LSPs). The selected protocols were chosen based on the value of staked ETH [10] and represent all previously described mechanisms. A summary of the design decisions for these LSPs is presented in Table V.
TABLE V: Taxonomy of Liquid Staking Protocols (LSPs) comparing permissionless (open) access to operate a validator, collateral requirements to run operate a validator and visibility of validator sets.
@@ IV-B Empirical Analysis
Table VI presents the taxonomy of the major LSTs that were empirically analyzed in this study, with most tokens operating in the reward-based model and naively launched on Ethereum. The differentiations between rebase and reward models for staking reward distribution is crucial for correct empirical analysis of the tokens.
TABLE VI: Token taxonomy for Liquid Staking Tokens (LSTs) with staking reward distribution models supported blockchains. The analyzed tokens are naively minted only on Ethereum and bridged to other blockchain.
@@
StakeWise sETH2 Dual Ethereum
rETH2
@@
The dispersion analysis of ETH LSTs is presented in Figure 7. This analysis highlights: (i) The distribution of daily return differences between reward-LSTs and daily staking rewards. ii) The deviation from the target value of 1 ETH for rebase-LSTs.
@@ V-E Comparison of Restaking Protocols
Restaking and liquid restaking represent emerging advancements in the DeFi space, garnering significant user attention and capital. Collectively, restaking protocols boast a TVL of over \$25.2 billion, with the majority concentrated in three major protocols: EigenLayer (\$15.07 billion), Babylon Protocol (\$5.5 billion), and Symbiotic (\$2.2 billion)[63]. Notably, EigenLayer is the third-largest DeFi protocol by TVL. A brief overview of these protocols is given in Table VII.
TABLE VII: Overview of major restaking protocols, highlighting total value locked (TVL), underlying assets, Actively Validated Services (AVS), Governance (Gov.) Token and associated metrics.
Protocol Reference TVL (\$B) Underlying Assets Restaking Rewards Gov. Token AVS